# Morita equivalence of Nijenhuis structures

### v1

@@ 4.3 Global and infinitesimal Morita equivalences of compatible structures
$\begin{cases}\mathrm{d}\mu_{i}\circ J=r_{i}\circ\mathrm{d}\mu_{i},\\ J\circ\pi_{\varpi}^{\sharp}\circ\mu_{i}^{*}=\pi_{\varpi}^{\sharp}\circ\mu_{i}^{*}\circ r_{i}^{*},\\ \nabla^{J}_{v}\left(\pi_{\varpi}^{\sharp}(\mu_{1}^{*}\alpha+\mu_{2}^{*}\beta)\right)=\pi_{\varpi}^{\sharp}\left(\mu_{1}^{*}\left(\nabla^{r_{1},*}_{\mathrm{d}\mu_{1}(v)}\alpha\right)+\mu_{2}^{*}\left(\nabla^{r_{2},*}_{\mathrm{d}\mu_{2}(v)}\beta\right)\right),\end{cases}$ (51)
@@
Holomorphic Dirac structures can be equivalently described in terms of the Nijenhuis structure $I_{M}$ of the complex manifold [8, Proposition 3.16], using the map
@@
Via this identification, we can see that two holomorphic Dirac structures, $L_{1}\subseteq\mathbb{T}^{(1,0)}M_{1}$ and $L_{2}\subseteq\mathbb{T}^{(1,0)}M_{2}$, are Morita equivalent if and only if there is a holomorphic closed $2$-form $\varpi\in\Omega^{(2,0)}(P)$, and two holomorphic forward Dirac maps $\left(M_{1},L_{1}\right)\xleftarrow{\ \mu_{1}\ }(P,\varpi)\xrightarrow{\ \mu_{2}\ }\left(M_{2},\bar{L}_{2}\right)$, satisfying (a) and (b) in Definition 4.11, and the actions $\mathscr{a}_{i}:\Gamma(L_{i})\to\mathfrak{X}(P)$ defined by $(\mathrm{d}\mu_{i}(v),\alpha)\mapsto v$ preserve the holomorphic structure. In particular, Morita equivalences of holomorphic Poisson structures [24] are described in this way.
@@
Let $(\mathscr{G},\omega_{1},\eta_{1},N_{1})$ and $(\mathscr{H},\omega_{2},\eta_{2},N_{2})$ be quasi-symplectic-Nijenhuis groupoids and consider a Morita equivalence given by a Morita bibundle:
@@
Then the associated twisted Dirac-Nijenhuis structures are Morita equivalent via:
$(G_{0},L_{1},r_{1})\xleftarrow{\quad\mu_{1}\quad}(P,\varpi,J)\xrightarrow{\quad\mu_{2}\quad}(H_{0},L_{2},r_{2}),$
with infinitesimal actions induced by the global actions.
@@
The associated infinitesimal Nijenhuis structures are Morita equivalent due to Theorem 3.18, and the compatibility of the pair $(\varpi,J)$ holds by hypothesis. Thus, it remains to show that the Dirac structures are Morita equivalent.
The quasi-symplectic Morita equivalence implies the existence of an action $\mathscr{A}:K_{1}\times_{(\mathbf{s},\mu)}P_{0}\to P_{0}$ of the quasi-symplectic groupoid $\mathscr{K}\coloneqq\mathscr{G}\times\mathscr{H}^{\mathrm{op}}$ (with $2$-form $\omega\coloneqq\mathrm{pr}_{G_{1}}^{*}\omega_{1}-\mathrm{pr}_{H_{1}}^{*}\omega_{2}$) on the submersion $\mu\coloneqq(\mu_{1}(p)\times\mu_{2}(p)$ satisfying
$\operatorname{ker}\mathrm{d}\mu\cap\operatorname{ker}\varpi=0\qquad\text{and}\qquad\mathscr{A}^{*}\varpi=\mathrm{pr}_{P_{0}}^{*}\varpi+\mathrm{pr}_{K_{1}}^{*}\omega.$
This implies that $P_{0}$ is a hamiltonian $\mathscr{K}$-space. Therefore, by [7, Theorem 4.7], the map $\mu:P_{0}\to G_{0}\times H_{0}$ is a forward Dirac map between $L_{\varpi}$ and the Dirac structure $L_{1}\times\bar{L}_{2}$ induced by $(\mathscr{K},\omega,\mathrm{pr}_{1}^{*}\eta_{1}-\mathrm{pr}_{2}^{*}\eta_{2})$.
Let $A_{\mathscr{G}}$ and $A_{\mathscr{H}}$ denote the Lie algebroids of $\mathscr{G}$ and $\mathscr{H}$, respectively, and let $(\rho_{1},\sigma_{\omega_{1}}):A_{\mathscr{G}}\to L_{1}$ and $(\rho_{2},\sigma_{\omega_{2}}):A_{\mathscr{H}}\to L_{2}$ be the isomorphisms of the Lie algebroids with the Dirac structures. Since the Lie groupoids are Morita equivalent, their corresponding Lie algebroids are Morita equivalent by Lemma 3.10. By Definition 3.8, we have that $\mathscr{a}_{1}(\mu^{*}_{1}A_{\mathscr{G}})_{p}=\operatorname{ker}\mathrm{d}\mu_{2\,p}$ and $\mathscr{a}_{2}(\mu_{2}^{*}A_{\mathscr{H}})_{p}=\operatorname{ker}\mathrm{d}\mu_{1\,p}$ for every $p\in P_{0}$. Consider $u\in\operatorname{ker}\mathrm{d}\mu_{1\,}$ and $v\in\operatorname{ker}\mathrm{d}\mu_{2\,p}$, and let $\xi\in A_{\mathscr{G}}$ and $\zeta\in A_{\mathscr{H}}$ be such that $\mathscr{a}_{2}(\zeta)=u$ and $\mathscr{a}_{1}(\xi)=v$. Since $\mu_{2}$ is a forward Dirac map we have in particular that $(\mathrm{d}\mu_{2})^{*}(\sigma_{\omega_{2}}(\zeta))=-\varpi^{\flat}(u)$, so
$\varpi(u,v)=\varpi^{\flat}(u)(v)=-(\mathrm{d}\mu_{2})^{*}\left(\sigma_{\omega_{2}}(\zeta)\right)(v)=-\sigma_{\omega_{2}}(\zeta)(\mathrm{d}\mu_{2}(v))=0.$
Therefore, condition (b) in Definition 4.11 is satified. As the remaining conditions hold by hypothesis, the Dirac structures are Morita equivalent.
@@
Consider two Dirac-Nijenhuis structures $(M_{1},L_{1},r_{1})$ and $(M_{2},L_{2},r_{2})$, where $L_{1}$ is $\eta_{1}$-twisted and $L_{2}$ is $\eta_{2}$-twisted, Morita equivalent via a bibundle:
$(M_{1},L_{1},r_{1})\xleftarrow{\quad\mu_{1}\quad}\left(P,\varpi,J\right)\xrightarrow{\quad\mu_{2}\quad}(M_{2},L_{2},r_{2}).$
Then the source-simply connected quasi-symplectic-Nijenhuis groupoids $(\mathscr{G},\omega_{1},\eta_{1},N_{1})$ and $(\mathscr{H},\omega_{2},\eta_{2},N_{2})$ integrating $(L_{1},r_{1})$ and $(L_{2},r_{2})$ are Morita equivalent via:
with global actions integrating the infinitesimal actions.
@@
Due to Proposition 4.13 the Lie algebroids of $L_{1}$ and $L_{2}$ are Morita equivalent Lie algebroids. By Lemma 3.10 this Morita equivalence is integrated to a Morita equivalence of the Lie groupoids $\mathscr{G}$ and $\mathscr{H}$. Consider the quasi-symplectic-Nijenhuis groupoids $(\mathscr{G},\omega_{1},N_{1})$ and $(\mathscr{H},\omega_{2},N_{2})$ integrating $(L_{1},r_{1})$ and $(L_{2},r_{2})$, respectively. By Theorem 3.19 we know that the underlying Nijenhuis groupoids are Morita equivalent, and the compatibility of the pair $(\varpi,J)$ is given by hypothesis. Therefore we just need to prove that the sum of the infinitesimal actions integrates to an action $\mathscr{A}$ that satisfies
$\mathscr{A}^{*}\varpi=\mathrm{pr}_{P_{0}}^{*}\varpi+\mathrm{pr}_{G_{1}}^{*}\omega_{1}-\mathrm{pr}_{H_{1}}^{*}\omega_{2}$
since all the other conditions of the quasi-symplectic Morita equivalence are given.
Note that the forward Dirac map $\mu:P_{0}\to M_{1}\times M_{2}$ between $L_{\varpi}$ and $L_{1}\times\bar{L}_{2}$ is a complete Dirac realization, i.e. the sum of the actions (Equation (30)) of $L_{1}$ and $L_{2}$ is complete. Indeed, since both actions $\mathscr{a}_{1}$ and $\mathscr{a}_{2}$ are complete, for every compactly supported sections $\xi\in\Gamma(L_{1})$ and $\zeta\in\Gamma(L_{2})$ the vector fields $\mathscr{a}_{1}(\xi)$ and $\mathscr{a}_{2}(\zeta)$ are complete, and by the Morita equivalence conditions they commute, therefore the sum $\mathscr{a}_{1}(\xi)+\mathscr{a}_{2}(\zeta)$ is complete. By [7, Theorem 4.7], the complete Dirac realization integrates to an action $\mathscr{A}$ of $\mathscr{K}=\mathscr{G}\times\mathscr{H}^{\mathrm{op}}$ on the submersion $\mu(p)=(\mu_{1}(p),\mu_{2}(p))$ such that
$\mathscr{A}^{*}\varpi=\mathrm{pr}_{P_{0}}^{*}\varpi+\mathrm{pr}_{K_{1}}^{*}\omega,$
where $\omega\in\Omega^{2}\left(\mathscr{K}\right)$ is the multiplicative $2$-form integrating $L_{1}\times\bar{L}_{2}$, i.e. $\omega=\mathrm{pr}_{1}^{*}\omega_{1}-\mathrm{pr}_{2}^{*}\omega_{2}$, which concludes the proof considering that $\mathrm{pr}_{G_{1}}\circ\mathrm{pr}_{K_{1}}=\mathrm{pr}_{G_{1}}$ and $\mathrm{pr}_{H_{1}}\circ\mathrm{pr}_{K_{1}}=\mathrm{pr}_{H_{1}}$. ∎
@@
4.4 Remarks on hierarchies of Morita equivalent Dirac-Nijenhuis structures
@@
Recall that for a given Poisson-Nijenhuis manifold $(M,\pi,r)$, the hierarchy of compatible structures is a sequence of Poisson bivectors $\{\pi^{n}\}_{n\in\mathbb{N}}$, defined recursively by
$\pi^{n\ \sharp}\coloneqq r\circ\pi^{(n-1)\ \sharp}=r^{n}\circ\pi^{\sharp}.$ (52)
This construction admits a generalization to the setting of Dirac-Nijenhuis structures [8]. Let $(L,r)$ be a Dirac-Nijenhuis structure on a manifold $M$. For every $n\in\mathbb{N}$, define:
$L^{(n,0)}\coloneqq\left(r^{n},\mathrm{id}_{T^{*}M}\right)(L)\qquad\text{and}$ (53)
$L^{(0,n)}\coloneqq\left(\mathrm{id}_{TM},(r^{*})^{n}\right)(L).$ (54)
According to [8, Proposition 4.8], if $\operatorname{ker}(r,\mathrm{id}_{T^{*}M})|_{L}=0$, then $(L^{(n,0)},r)$ is a Dirac-Nijenhuis structure for every $n\in\mathbb{N}$. Similarly, if $\operatorname{ker}(\mathrm{id}_{TM},r^{*})|_{L}=0$ holds, then $(L^{(0,n)},r)$ is a Dirac-Nijenhuis structure.
We now turn to the Morita equivalence of these hierarchies. Note that given a Morita equivalence $(M_{1},L_{1},r_{1})\leftarrow(P,\varpi,J)\to(M_{2},L_{2},r_{2})$, the Dirac structures $L_{\varpi}^{(0,n)}$ are graphs of $2$-forms; hence, they are natural candidates for bibundles between $L_{1}^{(0,n)}$ and $L_{2}^{(0,n)}$. In contrast, the structures $L_{\varpi}^{(n,0)}$ are not generally graphs of $2$-forms, so the hierarchies $L_{1}^{(n,0)}$ and $L_{2}^{(n,0)}$ given by (53) require separate analysis. Poisson-Nijenhuis structures yield hierarchies of the form (53), so the investigation of Morita equivalence in this context requires extra refinement. Thus, we proceed first to study the Morita equivalence of Dirac structures defined by the recursion (54). It turns out that, in this setting, the corresponding hierarchies are indeed Morita equivalent. Subsequently, we examine hierarchies of the form (53) specifically for the case of Poisson-Nijenhuis structures.
@@
Let $(M_{1},L_{1},r_{1})$ and $(M_{2},L_{2},r_{2})$ be Morita equivalent Dirac-Nijenhuis structures. If $\operatorname{ker}\left(\mathrm{id}_{TM_{i}},r_{i}^{*}\right)|_{L_{i}}=0$ for $i=1,2$, then the Dirac-Nijenhuis structures $\big(L_{1}^{(0,n)},r_{1}\big)$ and $\big(L_{2}^{(0,n)},r_{2}\big)$ are Morita equivalent for every $n\in\mathbb{N}$.
@@
We claim that if $(M_{1},L_{1},r_{1})\xleftarrow{\ \mu_{1}\ }(P,\varpi,J)\xrightarrow{\ \mu_{2}\ }(M_{2},L_{2},r_{2})$ is a Morita equivalence of the Dirac-Nijenhuis then
$\left(M_{1},L_{1}^{(0,n)},r_{1}\right)\xleftarrow{\quad\mu_{1}\quad}(P,\varpi_{n},J)\xrightarrow{\quad\mu_{2}\quad}\left(M_{2},L_{2}^{(0,n)},r_{2}\right)$

### v2

@@ 4.3. Global and infinitesimal Morita equivalences of compatible structures
(51) $\begin{cases}\mathrm{d}\mu_{i}\circ J=r_{i}\circ\mathrm{d}\mu_{i},\\ J\circ\pi_{\varpi}^{\sharp}\circ\mu_{i}^{*}=\pi_{\varpi}^{\sharp}\circ\mu_{i}^{*}\circ r_{i}^{*},\\ \nabla^{J}_{v}\left(\pi_{\varpi}^{\sharp}(\mu_{1}^{*}\alpha+\mu_{2}^{*}\beta)\right)=\pi_{\varpi}^{\sharp}\left(\mu_{1}^{*}\left(\nabla^{r_{1},*}_{\mathrm{d}\mu_{1}(v)}\alpha\right)+\mu_{2}^{*}\left(\nabla^{r_{2},*}_{\mathrm{d}\mu_{2}(v)}\beta\right)\right),\end{cases}$
@@
Holomorphic Dirac structures can be equivalently described in terms of the Nijenhuis structure $I_{M}$ of the complex manifold [undefh, Proposition 3.16], using the map
@@
Via this identification, we can see that two holomorphic Dirac structures, $L_{1}\subseteq\mathbb{T}^{(1,0)}M_{1}$ and $L_{2}\subseteq\mathbb{T}^{(1,0)}M_{2}$, are Morita equivalent if and only if there is a holomorphic closed 22-form $\varpi\in\Omega^{(2,0)}(P)$, and two holomorphic forward Dirac maps $\left(M_{1},L_{1}\right)\xleftarrow{\ \mu_{1}\ }(P,\varpi)\xrightarrow{\ \mu_{2}\ }\left(M_{2},\bar{L}_{2}\right)$, satisfying (a) and (b) in Definition 4.11, and the actions $\mathscr{a}_{i}:\Gamma(L_{i})\to\mathfrak{X}(P)$ defined by $(\mathrm{d}\mu_{i}(v),\alpha)\mapsto v$ preserve the holomorphic structure. In particular, Morita equivalences of holomorphic Poisson structures [undefx] are described in this way.
@@
Let $(\mathscr{G},\omega_{1},\eta_{1},N_{1})$ and $(\mathscr{H},\omega_{2},\eta_{2},N_{2})$ be quasi-symplectic-Nijenhuis groupoids. Assume they are Morita equivalent via a Morita $(\mathscr{G},\mathscr{H})$-bibundle
@@
Then their associated twisted Dirac-Nijenhuis structures, $(L_{1},r_{1})$ and $(L_{2},r_{2})$, are Morita equivalent, with $(P,\varpi,J)$ serving as the infinitesimal Morita $(L_{1},L_{2})$-bibundle equipped with infinitesimal actions induced by the global actions.
@@
By Theorem 3.19, the associated infinitesimal Nijenhuis structures are Morita equivalent. Furthermore, the compatibility of the pair $(\varpi,J)$ holds by hypothesis. Therefore, it remains only to establish the Morita equivalence of the underlying Dirac structures.
The quasi-symplectic Morita equivalence yields an action $\mathscr{A}:K_{1}\times_{(\mathbf{s},\mu)}P\to P$ of the product quasi-symplectic groupoid $\mathscr{K}\coloneqq\mathscr{G}\times\mathscr{H}^{\mathrm{op}}$ on the submersion $\mu\coloneqq(\mu_{1},\mu_{2})$. Here, the 22-form on $\mathscr{K}$ is given by $\omega\coloneqq\mathrm{pr}_{G_{1}}^{*}\omega_{1}-\mathrm{pr}_{H_{1}}^{*}\omega_{2}$, and the action satisfies the conditions
$\operatorname{ker}\mathrm{d}\mu\cap\operatorname{ker}\varpi=0\qquad\text{and}\qquad\mathscr{A}^{*}\varpi=\mathrm{pr}_{P}^{*}\varpi+\mathrm{pr}_{K_{1}}^{*}\omega.$
Consequently, $P$ inherits the structure of a Hamiltonian $\mathscr{K}$-space. Applying [undeff, Theorem 4.7], we conclude that $\mu:P\to G_{0}\times H_{0}$ is a forward Dirac map from $L_{\varpi}$ to the product Dirac structure $L_{1}\times\bar{L}_{2}$ induced by $(\mathscr{K},\omega,\mathrm{pr}_{1}^{*}\eta_{1}-\mathrm{pr}_{2}^{*}\eta_{2})$.
Let $A_{\mathscr{G}}$ and $A_{\mathscr{H}}$ denote the Lie algebroids of $\mathscr{G}$ and $\mathscr{H}$, respectively. These Lie algebroids are isomorphic to their associated Dirac structures via the bundle maps $(\rho_{1},\sigma_{\omega_{1}}):A_{\mathscr{G}}\xrightarrow{\sim}L_{1}$ and $(\rho_{2},\sigma_{\omega_{2}}):A_{\mathscr{H}}\xrightarrow{\sim}L_{2}$. Because the underlying Lie groupoids are Morita equivalent, Lemma 3.11 ensures that their corresponding Lie algebroids are also Morita equivalent. By Definition 3.8, the associated infinitesimal actions must satisfy $\mathscr{a}_{1}(\mu^{*}_{1}A_{\mathscr{G}})_{p}=\operatorname{ker}\mathrm{d}\mu_{2\,p}$ and $\mathscr{a}_{2}(\mu_{2}^{*}A_{\mathscr{H}})_{p}=\operatorname{ker}\mathrm{d}\mu_{1\,p}$ for every $p\in P$. Therefore, for any tangent vectors $u\in\operatorname{ker}\mathrm{d}\mu_{1\,p}$ and $v\in\operatorname{ker}\mathrm{d}\mu_{2\,p}$ there exist elements $\mu_{1}^{*}\xi\in\mu_{1}^{*}A_{\mathscr{G}}$ and $\mu_{2}^{*}\zeta\in\mu_{2}^{*}A_{\mathscr{H}}$ such that $\mathscr{a}_{2}(\mu_{2}^{*}\zeta)=u$ and $\mathscr{a}_{1}(\mu_{1}^{*}\xi)=v$. Because $\mu_{2}$ is a forward Dirac map, it follows that $(\mathrm{d}\mu_{2})^{*}(\sigma_{\omega_{2}}(\zeta))=-\varpi^{\flat}(u)$. Evaluating $\varpi$ on the pair $(u,v)$ thus yields
$\varpi(u,v)=\varpi^{\flat}(u)(v)=-(\mathrm{d}\mu_{2})^{*}\left(\sigma_{\omega_{2}}(\zeta)\right)(v)=-\sigma_{\omega_{2}}(\zeta)(\mathrm{d}\mu_{2}(v))=0,$
where the final equality holds because $v\in\operatorname{ker}\mathrm{d}\mu_{2\,p}$. This confirms that the orthogonality condition (d) of Definition 4.11 is satisfied. Because the remaining conditions hold by hypothesis, the Morita equivalence of the Dirac structures is established.
@@
Let $\mathscr{G}=(G_{1}\rightrightarrows G_{0})$ and $\mathscr{H}=(H_{1}\rightrightarrows H_{0})$ be source-simply connected Lie groupoids. Suppose that the base spaces $G_{0}$ and $H_{0}$ admit Dirac–Nijenhuis structures $(G_{0},L_{1},r_{1})$ and $(H_{0},L_{2},r_{2})$, respectively, where $L_{1}$ is $\eta_{1}$-twisted and $L_{2}$ is $\eta_{2}$-twisted. Assume further that these structures are Morita equivalent via an infinitesimal Morita $(L_{1},L_{2})$-bibundle $(P,\varpi,J)$:
$(G_{0},L_{1},r_{1})\xleftarrow{\quad\mu_{1}\quad}(P,\varpi,J)\xrightarrow{\quad\mu_{2}\quad}(H_{0},L_{2},r_{2}).$
Then the corresponding quasi-symplectic–Nijenhuis groupoids $(\mathscr{G},\omega_{1},\eta_{1},N_{1})$ and $(\mathscr{H},\omega_{2},\eta_{2},N_{2})$ are Morita equivalent. In particular, $(P,\varpi,J)$ serves as the Morita $(\mathscr{G},\mathscr{H})$-bibundle, equipped with global actions that integrate the infinitesimal actions.
@@
Theorem 3.20 guarantees that the underlying Morita equivalence of the infinitesimal Nijenhuis structures integrates to yield Morita equivalent Nijenhuis groupoids.
Furthermore, the compatibility of the pair $(\varpi,J)$ holds by hypothesis. Because all other requisites for quasi-symplectic Morita equivalence are already established, it remains only to prove that the total infinitesimal action integrates to a global action $\mathscr{A}$ satisfying
$\mathscr{A}^{*}\varpi=\mathrm{pr}_{P}^{*}\varpi+\mathrm{pr}_{G_{1}}^{*}\omega_{1}+\mathrm{pr}_{H_{1}}^{*}\omega_{2}.$
Observe that the forward Dirac map $\mu:P\to G_{0}\times H_{0}$ from $L_{\varpi}$ to $L_{1}\times\bar{L}_{2}$ is a complete Dirac realization; that is, the infinitesimal action of $L_{1}\times\bar{L}_{2}$ on $P$ is complete. To verify this, note that because the individual actions $\mathscr{a}_{1}$ and $\mathscr{a}_{2}$ are complete, the induced vector fields $\mathscr{a}_{1}(\xi)$ and $\mathscr{a}_{2}(\zeta)$ are complete for any compactly supported sections $\xi\in\Gamma(L_{1})$ and $\zeta\in\Gamma(L_{2})$. Since these vector fields commute, their sum $\mathscr{a}_{1}(\xi)+\mathscr{a}_{2}(\zeta)$ is also complete, which establishes the completeness of the total infinitesimal action.
Applying [undeff, Theorem 4.7], this complete Dirac realization integrates to a global left action $\widetilde{\mathscr{A}}$ of the product groupoid $\mathscr{K}=\mathscr{G}\times\mathscr{H}^{\mathrm{op}}$ on the submersion $\mu=(\mu_{1},\mu_{2})$ satisfying $\widetilde{\mathscr{A}}^{*}\varpi=\mathrm{pr}_{P}^{*}\varpi+\mathrm{pr}_{K_{1}}^{*}\omega$, where $\omega\in\Omega^{2}(\mathscr{K})$ is the multiplicative 22-form integrating $L_{1}\times\bar{L}_{2}$, given explicitly by $\omega=\mathrm{pr}_{1}^{*}\omega_{1}-\mathrm{pr}_{2}^{*}\omega_{2}$. Applying the projection identities $\mathrm{pr}_{G_{1}}\circ\mathrm{pr}_{K_{1}}=\mathrm{pr}_{G_{1}}$ and $\mathrm{pr}_{H_{1}}\circ\mathrm{pr}_{K_{1}}=\mathrm{pr}_{H_{1}}$, the pullback condition expands to $\widetilde{\mathscr{A}}^{*}\varpi=\mathrm{pr}_{P}^{*}\varpi+\mathrm{pr}_{G_{1}}^{*}\omega_{1}-\mathrm{pr}_{H_{1}}^{*}\omega_{2}$.
The existence of this left $\mathscr{K}$-action $\widetilde{\mathscr{A}}$ corresponds precisely to equipping $P$ with commuting left Hamiltonian $\mathscr{G}$- and right Hamiltonian $\mathscr{H}$-actions. Let $\mathscr{A}$ denote the combination of this left $\mathscr{G}$-action and right $\mathscr{H}$-action. By reversing the sign of $\omega_{2}$—as consequence of converting the left $\mathscr{H}^{\mathrm{op}}$-action into a right $\mathscr{H}$-action—the combined action satisfies the desired compatibility condition $\mathscr{A}^{*}\varpi=\mathrm{pr}_{P}^{*}\varpi+\mathrm{pr}_{G_{1}}^{*}\omega_{1}+\mathrm{pr}_{H_{1}}^{*}\omega_{2}$. This concludes the proof. ∎
@@
$4.$ Remarks on hierarchies of Morita equivalent Dirac-Nijenhuis structures
@@
Recall that for a given Poisson-Nijenhuis manifold $(M,\pi,r)$, the hierarchy of compatible structures is a sequence of Poisson bivectors $\{\pi^{(n)}\}_{n\in\mathbb{N}}$, defined recursively by
(52) $\big(\pi^{(n)}\big)^{\sharp}\coloneqq r\circ\big(\pi^{(n-1)}\big)^{\sharp}=r^{n}\circ\pi^{\sharp}.$
This construction admits a generalization to the setting of Dirac-Nijenhuis structures [undefh]. Let $(L,r)$ be a Dirac-Nijenhuis structure on a manifold $M$. For every $n\in\mathbb{N}$, define:
(53) $L^{(n,0)}\coloneqq\left(r^{n},\mathrm{id}_{T^{*}M}\right)(L)\qquad\text{and}$
(54) $L^{(0,n)}\coloneqq\left(\mathrm{id}_{TM},(r^{*})^{n}\right)(L).$
According to [undefh, Proposition 4.8], if $\operatorname{ker}(r,\mathrm{id}_{T^{*}M})|_{L}=0$, then $(L^{(n,0)},r)$ is a Dirac-Nijenhuis structure for every $n\in\mathbb{N}$. Similarly, if $\operatorname{ker}(\mathrm{id}_{TM},r^{*})|_{L}=0$ holds, then $(L^{(0,n)},r)$ is a Dirac-Nijenhuis structure.
We now turn to the Morita equivalence of these hierarchies. Note that given a Morita equivalence $(M_{1},L_{1},r_{1})\leftarrow(P,\varpi,J)\to(M_{2},L_{2},r_{2})$, the Dirac structures $L_{\varpi}^{(0,n)}$ are graphs of 22-forms; hence, they are natural candidates for bibundles between $L_{1}^{(0,n)}$ and $L_{2}^{(0,n)}$. In contrast, the structures $L_{\varpi}^{(n,0)}$ are not generally graphs of 22-forms, so the hierarchies $L_{1}^{(n,0)}$ and $L_{2}^{(n,0)}$ given by (53) require separate analysis. Poisson-Nijenhuis structures yield hierarchies of the form (53), so the investigation of Morita equivalence in this context requires extra refinement. Thus, we proceed first to study the Morita equivalence of Dirac structures defined by the recursion (54). It turns out that, in this setting, the corresponding hierarchies are indeed Morita equivalent. Subsequently, we examine hierarchies of the form (53) specifically for the case of Poisson-Nijenhuis structures.
@@
Let $(M_{1},L_{1},r_{1})$ and $(M_{2},L_{2},r_{2})$ be Morita equivalent Dirac-Nijenhuis structures. Suppose that $\operatorname{ker}\left(\mathrm{id}_{TM_{i}},r_{i}^{*}\right)|_{L_{i}}=0$ for $i=1,2$. Then the Dirac-Nijenhuis structures $\big(M_{1},L_{1}^{(0,n)},r_{1}\big)$ and $\big(M_{2},L_{2}^{(0,n)},r_{2}\big)$ are Morita equivalent for every $n\in\mathbb{N}$.
@@
We claim that if $(M_{1},L_{1},r_{1})\xleftarrow{\ \mu_{1}\ }(P,\varpi,J)\xrightarrow{\ \mu_{2}\ }(M_{2},L_{2},r_{2})$ is a Morita equivalence of Dirac-Nijenhuis structures, then
$\left(M_{1},L_{1}^{(0,n)},r_{1}\right)\xleftarrow{\quad\mu_{1}\quad}\big(P,\varpi^{(n)},J\big)\xrightarrow{\quad\mu_{2}\quad}\left(M_{2},L_{2}^{(0,n)},r_{2}\right)$